OpenStudy (gabylovesyou):

Let f(x) = 2x - 6. Solve f^-1(x) when x = 2.

OpenStudy (gabylovesyou):

@hartnn

OpenStudy (gabylovesyou):

@TuringTest

OpenStudy (sweetburger):

Ok first off you want to say that F(x) = y so you write the equation y=2x-6. Now replace each x with a y and each y with a x like this x= 2y-6 now you would solve fore y like this y=(x+6)/2 then plug in x=2 for each x value like this y = (2+6)/2 which equates to y=4